OpenStudy (anonymous):

what is the equation of the line that passes through the points 4 , -6 and has a slope of -3

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

This can be solved by using the slope-intercept form, y=mx+b. Since you have the point (4, -6) this can be plugged in along with the slope -3. The x variable ends up being 4, the y variable ends up being -6, and the m variable proves to be -3 (m=slope). Now that you know this you can plug the variables into the equation to find b. so: (-6)=(-3)(4)+b (-6)=(-12)+b 6=b Therefore, the answer ends up to be: y=-3x+6. Hope this helped!